OpenStudy (anonymous):

(7x)^2/3 write each expression in radical form? please help

OpenStudy (anonymous):

\[\sqrt[3]{(7x)^2}\]

OpenStudy (anonymous):

With a fraction as an exponent, the numerator is the "radical" raised to a power. The denominator is the root of the radical. This is \[(\sqrt[3]{7x})^2\]

OpenStudy (anonymous):

Make sure you take the cube root of 7x BEFORE you raise the answer to the 2nd power. Remember the order of operations plays here...hence the parenthesis.

OpenStudy (anonymous):

so (6n)^5/2 is \[\sqrt[2]{6n^5}\]

OpenStudy (anonymous):

Almost. \[(\sqrt[2]{6n})^5\] The parenthesis matter. The exponent must be outside of the actual radical.